# History, Development, and Principles of Large Language Models—An Introductory Survey

### v2

@@ 5. DRAWBACKS AND FUTURE DIRECTIONS OF LARGE LANGUAGE MODELS
Fairness: LLMs are widely employed for decision-making across various domains, impacting individuals and society. However, these decisions might exhibit bias against marginalized populations. For instance, when utilized in screening resumes for programming positions, LLMs might exhibit a bias favoring male candidates, demonstrating evident gender discrimination. This bias stems from the training of LLMs on extensive and unstructured Internet data, where they inherit stereotypes, misrepresentations, derogatory language, and exclusionary behaviors. These aspects disproportionately affect already disadvantaged and marginalized groups [bender2021dangers, dodge2021documenting, sheng2021societal]. The commonly employed strategy to mitigate bias in LLMs is to remove biased data from the training dataset. However, this strategy does not entirely eliminate bias and often diminishes the effectiveness of language modeling [meade2021empirical]. [saxena2023missed] discuss that the root causes of these problems lie not only in technology but also in socially acceptable definitions of fairness and meaningful interventions to ensure the long-term well-being of all groups. Researchers should consider the needs of disadvantaged groups from the outset and design these technologies proactively, rather than simply reacting to the biases present in their designed systems [gallegos2023bias]. The documented bias in LLMs, as highlighted by studies like [blodgett2017racial, mei2023bias, mozafari2020hate], underscores the imperative for future efforts in this direction. Fairness: LLMs are widely employed for decision-making across various domains, impacting individuals and society. However, these decisions might exhibit bias against marginalized populations. For instance, when utilized in screening resumes for programming positions, LLMs might exhibit a bias favoring male candidates, demonstrating evident gender discrimination. This bias stems from the training of LLMs on extensive and unstructured Internet data, where they inherit stereotypes, misrepresentations, derogatory language, and exclusionary behaviors. These aspects disproportionately affect already disadvantaged and marginalized groups [bender2021dangers, dodge2021documenting, sheng2021societal]. The commonly employed strategy to mitigate bias in LLMs is to remove biased data from the training dataset. However, this strategy does not entirely eliminate bias and often diminishes the effectiveness of language modeling [meade2021empirical]. [saxena2023missed] discuss that the root causes of these problems lie not only in technology but also in socially acceptable definitions of fairness and meaningful interventions to ensure the long-term well-being of all groups. Researchers should consider the needs of disadvantaged groups from the outset and design these technologies proactively, rather than simply reacting to the biases present in their designed systems [gallegos2023bias]. The documented bias in LLMs, as highlighted by studies like [blodgett2017racial, mei2023bias], underscores the imperative for future efforts in this direction.
@@
Safety: LLMs have diverse applications in industries such as finance and healthcare. However, these applications also present significant security challenges. LLMs may generate outputs that diverge from the provided context, user input, or factual knowledge, a phenomenon referred to as “hallucination” [bang2023multitask, huang2023survey]. This undermines their reliability in real-world scenarios, emphasizing the vital necessity for precision, particularly in critical domains like medicine [dash2023evaluation, umapathi2023med]. As a primary way to avoid these problems, well-aligned LLMs can be developed by including humans in the training loop and using reinforcement learning from human feedback (RLHF) [ouyang2022training, gao2020pile]. To improve model safety, it is also important to include safety-related cues in the RLHF process, as shown in GPT-4 [achiam2023gpt]. However, RLHF relies heavily on high-quality human feedback data from professional annotators, which is costly and time-consuming. As a result, enhancing the RLHF framework is essential to alleviate the burden on human annotators. Additionally, exploring more efficient annotation methods with assured data quality, such as Large Language Models (LLMs), can assist in the annotation process and enhance overall safety.
@@
Intellectual Property: LLMs possess the capability to create content resembling human-generated material, potentially encroaching upon users’ intellectual property rights. These models rely on training and optimizing using users’ textual data, which may encompass personal information, proprietary knowledge, patented technology, and more. Notably, AI-generated code exemplifies this issue. Code-generation models trained on open-source repositories can produce programs resembling existing ones, potentially disregarding relevant licenses [yu2023codeipprompt]. For instance, legal action has been taken against Microsoft, GitHub, and OpenAI, alleging copyright infringement due to Copilot reproducing licensed code without adherence to licensing terms [^11] https://www.reuters.com/technology/microsoft-attracting-users-its-code-writing-generative-ai-software-2023-01-25/  [^12] https://www.reuters.com/legal/litigation/openai-microsoft-want-court-toss-lawsuit-accusing-them-abusing-open-source-code-2023-01-27/. Consequently, there’s a pressing need for heightened attention and regulation concerning intellectual property rights concerning LLMs. This is essential to safeguard the legitimate rights and interests of original creators, users, and the public while fostering and respecting the innovation and advancement of LLMs technology [li2023protecting].
@@ 6. CONCLUSIONS
In recent times, Large Language Models (LLMs) have garnered significant attention from the fields of science, industry, society, and beyond. However, despite the substantial advantages that LLMs offer in enhancing both professional and personal aspects of life, a limited understanding among general practitioners regarding the background and principles of these models impedes their full potential. In light of this, the survey at hand systematically delves into the evolution of LLMs, presenting key concepts and techniques for a comprehensive understanding of these models. Moreover, the survey sheds light on the limitations of LLMs and identifies promising directions for future research. Notably, setting itself apart from other surveys on LLMs, this study prioritizes empowering practitioners, irrespective of their background knowledge, to ensure their proficient utilization in both scientific research and daily tasks.
CONFLICT OF INTEREST
[…]

### v3

@@ 5. DRAWBACKS AND FUTURE DIRECTIONS OF LARGE LANGUAGE MODELS
Privacy: LLMs are widely used across various sectors, but they also present significant privacy concerns [145, 65]. One major issue is the risk of data leakage, where LLMs may inadvertently memorize and reproduce sensitive or personal information from their training datasets, potentially leading to privacy breaches. This risk is particularly severe in domains like healthcare and finance, where exposing private data can have serious consequences [146]. Additionally, the training datasets used for LLMs are often not fully anonymized, meaning personal identifiers and other sensitive information could be embedded within the model [147]. This increases the risk of violating privacy regulations, such as the General Data Protection Regulation [148]. Recent developments have further complicated the privacy landscape, as tools like FraudGPT and WormGPT have emerged. These tools, designed for cybercrime, highlight the misuse potential of LLMs by generating fraudulent emails, suggesting malicious link placements, and facilitating cyberattacks. FraudGPT [149, 150] and WormGPT [151] were trained on datasets focused on malware and fraud, enabling cybercriminals to conduct sophisticated attacks such as Business Email Compromise (BEC). This illustrates the urgent need for robust privacy safeguards and regulatory frameworks to prevent the misuse of LLMs and protect sensitive information from being exploited maliciously.
@@
Fairness: LLMs are widely employed for decision-making across various domains, impacting individuals and society. However, these decisions might exhibit bias against marginalized populations [152, 153, 154, 155, 156]. For instance, when utilized in screening resumes for programming positions, LLMs might exhibit a bias favoring male candidates, demonstrating evident gender discrimination [157, 158, 159, 160, 161]. This bias stems from the training of LLMs on extensive and unstructured Internet data, where they inherit stereotypes, misrepresentations, derogatory language, and exclusionary behaviors. These aspects disproportionately affect already disadvantaged and marginalized groups [162, 15, 22]. The commonly employed strategy to mitigate bias in LLMs is to remove biased data from the training dataset. However, this strategy does not entirely eliminate bias and often diminishes the effectiveness of language modeling [163]. Saxena et al. [152] discussed that the root causes of these problems lie not only in technology but also in socially acceptable definitions of fairness and meaningful interventions to ensure the long-term well-being of all groups. Researchers should consider the needs of disadvantaged groups from the outset and design these technologies proactively, rather than simply reacting to the biases present in their designed systems [40, 164, 165]. The documented bias in LLMs, as highlighted by studies like [166, 167, 54], underscores the imperative for future efforts in this direction.
@@
Safety: LLMs have diverse applications in industries such as finance and healthcare. However, these applications also present significant security challenges. LLMs may generate outputs that diverge from the provided context, user input, or factual knowledge, a phenomenon referred to as “hallucination” [21, 66]. This undermines their reliability in real-world scenarios, emphasizing the vital necessity for precision, particularly in critical domains like medicine [168, 169]. As a primary way to avoid these problems, well-aligned LLMs can be developed by including humans in the training loop and using reinforcement learning from human feedback (RLHF) [109, 75]. To improve model safety, it is also important to include safety-related cues in the RLHF process, as shown in GPT-4 [11]. However, RLHF relies heavily on high-quality human feedback data from professional annotators, which is costly and time-consuming. As a result, enhancing the RLHF framework is essential to alleviate the burden on human annotators. Additionally, exploring more efficient annotation methods with assured data quality, such as LLMs, can assist in the annotation process and enhance overall safety.
@@
Intellectual Property: LLMs possess the capability to create content resembling human-generated material, potentially encroaching upon users’ intellectual property rights. These models rely on training and optimizing using users’ textual data, which may encompass personal information, proprietary knowledge, patented technology, and more. Notably, AI-generated code exemplifies this issue. Code-generation models trained on open-source repositories can produce programs resembling existing ones, potentially disregarding relevant licenses [34]. For instance, legal action has been taken against Microsoft, GitHub, and OpenAI, alleging copyright infringement due to Copilot reproducing licensed code without adherence to licensing terms  [^12] https://www.reuters.com/technology/microsoft-attracting-users-its-code-writing-generative-ai-software-2023-01-25/  [^13] https://www.reuters.com/legal/litigation/openai-microsoft-want-court-toss-lawsuit-accusing-them-abusing-open-source-code-2023-01-27/. Additionally, there are concerns about potential copyright infringements with LLM-generated images and texts [170, 171], as highlighted in recent lawsuits by a group of novelists against OpenAI in July 2023 for allegedly using their books without permission to train models [172], and by The New York Times against OpenAI and Microsoft in December 2023 for scraping articles to train their generative models without consent [173]. Consequently, there’s a pressing need for heightened attention and regulation concerning intellectual property rights concerning LLMs. This is essential to safeguard the legitimate rights and interests of original creators, users, and the public while fostering and respecting the innovation and advancement of LLMs technology [174].
@@ 6. CONCLUSIONS
In recent times, Large Language Models (LLMs) have garnered significant attention from the fields of science, industry, society, and beyond. However, despite the substantial advantages that LLMs offer in enhancing both professional and personal aspects of life, a limited understanding among general practitioners regarding the background and principles of these models impedes their full potential. In light of this, the survey at hand systematically delves into the evolution of LLMs, presenting key concepts and techniques for a comprehensive understanding of these models. It introduces a detailed taxonomy of LLMs and compares state-of-the-art models to assist readers in selecting the most suitable model for their specific needs. Additionally, the survey highlights the limitations of current LLMs and identifies promising areas for future research. Notably, setting itself apart from other surveys on LLMs, this study prioritizes empowering practitioners, irrespective of their background knowledge, to ensure their proficient utilization in both scientific research and daily tasks.
Acknowledgement
This work was supported in part by the National Science Foundation (NSF) under Grant No. 2245895.